OpenStudy (anonymous):

-2x+x=l5-9l How do i solve this and what do the bars mean?

OpenStudy (anonymous):

the bars mean absolute value

OpenStudy (saifoo.khan):

Those are called MODS or Modulus. What they do, they convert a negative number to a positive one. Let's say: -5. |-5| = 5

OpenStudy (anonymous):

so it would= -5+9?

OpenStudy (anonymous):

or 5+9?

OpenStudy (anonymous):

it would be 5-9 but still in the bars when you subtract 5 from 9 it becomes -4 and it is still in the bars therefore it becomes a positive 4 your equation would now be -2x+x=4

OpenStudy (saifoo.khan):

-2x+x=l5-9l -2x + x = |-4| -2x + x = +4

OpenStudy (anonymous):